OpenStudy (anonymous):

i need help with finding domain and range in functions give me time to write the equation

OpenStudy (ash2326):

sure

OpenStudy (anonymous):

\[\sqrt[1/2]{x?}\]

OpenStudy (anonymous):

w/o the ?

OpenStudy (anonymous):

alright but 1/2 is exponent ???

OpenStudy (ash2326):

Is this your question? \[\sqrt x \]

OpenStudy (anonymous):

no it has a 1/2 in front of it juss like i put but its supposed to be beside it not in it like tht

OpenStudy (anonymous):

\[1/2\sqrt{x}\]

OpenStudy (anonymous):

okay is okay

OpenStudy (anonymous):

thnks lol sorry

OpenStudy (ash2326):

oh you mean \[\frac 12 \sqrt x\] ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

domain si from 0 to infintiy range is from 0 to infintiy

OpenStudy (ash2326):

You know in a radical there can be only positive numbers, so domain is??

OpenStudy (ash2326):

@moha_10 Please don't just give out answers, please explain how you got them?

OpenStudy (anonymous):

no she has really been helping me she don't juss give out answers

OpenStudy (anonymous):

@ash2326 she has helped me a whole lot and she explains very well thnks

OpenStudy (ash2326):

That's good, did you understand how the domain and range are found?

OpenStudy (anonymous):

yes